OpenStudy (anonymous):

Solve the following equation. Please show work -8/13 x=24

OpenStudy (anonymous):

is this \[-\frac{8}{13}x\]

OpenStudy (anonymous):

or \[-\frac{8}{13x}\] put parentheses.

OpenStudy (anonymous):

its - 8 /13 x =24

OpenStudy (anonymous):

where are the parentheses? (-8/13)x=24 or what? -8/(13x)=24